Kyutiedask10:

Given the function f(x) = 2x, find the value of f−1(16). f−1(16) = 4 f−1(16) = 8 f−1(16) = 32 f−1(16) = 65,536

Ultrilliam:

what do you think it is?

Ultrilliam:

f-1 (16) = 8

Ultrilliam:

._.

Kyutiedask10:

I thought it was C

Ultrilliam:

that would be f−1(32)

Kyutiedask10:

or A

Ultrilliam:

that would be f−1(8)

Ultrilliam:

have you noticed the pattern yet? :P

Kyutiedask10:

nope

Ultrilliam:

in this instance f−1(x) x/2 = f−1(x)

Ultrilliam:

16/2=8 so f−1(16) = 8

Kyutiedask10:

ok I think I got it.. Thanks

Ultrilliam:

np :)